# 3D Question Answering via only 2D Vision-Language Models

### v1

@@ 5.1 Comparisons with the State-of-the-Arts
Table 1 presents the quantitative results comparing 2D-only methods (uniform sampling, image retrieval, and cdViews) with other LLAVA-OV (Li et al., 2024a) with state-of-the-art 3D and hybrid methods. First, it is observed that 2D-only methods achieve superior performance, showing the advantage of applying 2D pre-trained models for 3D tasks. For example, compared to BridgeQA (Mo & Liu, 2024), our $\mathcal{F}_{\texttt{cdViews}}$ achieves significant improvements of 15.0% and 21.1% CIDEr on the two test sets of ScanQA. Second, among the 2D-only methods, cdViews outperforms the others. For example, $\mathcal{F}_{\texttt{cdViews}}$ outperforms $\mathcal{F}_{\text{retrieval}}$ by 4.0% and 2.7% CIDEr on both test sets of ScanQA. The reason is that the uniform sampling method ignores the question and the image retrieval method often fails to capture critical views or introduces redundancy views. In contrast, cdViews effectively identifies critical and diverse views for efficient 3D-QA. The qualitative comparison of selected views is shown in Figure 2 and Figure 6. More comparisons are provided in the Appendix Section B.
@@
+ $\mathcal{F}_{\text{uniform}}$ 28.3 17
+ $\mathcal{F}_{retrieval}$ 29.1 17
+ $\mathcal{F}_{retrieval}$ 29.2 9
+ $\mathcal{F}_{\texttt{cdViews}}$ 29.7 17
@@ 5.2 Ablation Study
In this section, we conduct an ablation study on the validation set of ScanQA (Azuma et al., 2022), following (Mo & Liu, 2024). We study the impact of cdViews components and viewNMS thresholds. In addition, we particularly compare ours with the most related work: image-retrieval-based 3D-QA (Mo & Liu, 2024). More ablation studies are in Section C of the Appendix.
@@
Figure 7: The results of EM@1 using two configurations: optimal $k$ (blue) vs. fixed $k\!=\!9$ (green). X-axis is the threshold $T$ of viewNMS. $T\!=\!0$ means disabling viewNMS.
@@
Parameters $644$M $5.9$M ($-99.1\%$)
FLOPs $593.6$T $294.5$T ($-50.4\%$)
Inference Time $2.8$s $1.2$s ($-57.1\%$)
@@ Appendix A More Details in View Matching
This supplementary is for Sec. 4.1 of the main paper. In our viewAnnotator, view matching classifies views as positive, negative, or uncertain. However, directly using a 2D LVLM with the prompt $Prompt_{M}$ as an instruction is unreliable, as the model lacks an explicit judgment criterion. To address this, we leverage its strong in-context learning ability (Zhou et al., 2024) by providing a textual context example that guides the model through a structured reasoning process. Specifically, we incorporate a step-by-step system prompt in the View Matching process. As shown in Figure S1, the system prompt ensures that all key objects, attributes, and spatial relationships in the caption align with the image, reducing ambiguity and improving consistency. Uncertain views are explicitly excluded, enhancing the robustness of the annotation process. Additional examples of positive and negative views are shown in Figure S2.
@@
            ScanQA (Azuma et al., 2022) 3D 20.3 29.5 32.4 61.7
            3D-LLM (Hong et al., 2023) 3D 20.5 39.3 35.7 69.4
            3D-VLP (Jin et al., 2023a) 3D 21.7 30.5 34.5 67.0
            LL3DA (Chen et al., 2024a) 3D+2D - - 37.3 76.8
            BridgeQA (Mo & Liu, 2024) 3D+2D 27.0 - - -
            GPT-4O+CC (Liu et al., 2024a) 2D - 35.4 42.6 87.0
@@
margin over the compared best $3.1\uparrow$ $3.3\uparrow$ $4.2\uparrow$ $7.0\uparrow$
Table S1: Result comparisons with the state-of-the-art methods on the validation set of ScanQA (Azuma et al., 2022). The best and second best results are in bold and underlined.
@@ Appendix B More Comparisons with the State-of-the-Art Methods
This supplementary is for Section 5.1 of the main paper. Table S1 presents the quantitative results comparing LLAVA-OV (Li et al., 2024a) with different view selection methods, including uniform sampling, image retrieval, and our cdViews, against state-of-the-art methods on the validation set of ScanQA. As shown, LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ outperforms these methods by clear margins. The last row of Table 1 highlights the performance gap between LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ and the best-performing baselines. Even compared to GPT-4O+CC (Liu et al., 2024a), which leverages the powerful capabilities of GPT-4O (OpenAI, 2024), LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ surpasses it by 7.0% CIDEr. GPT-4O+CC improves spatial understanding by adding object markers to track correspondences across uniformly sampled views. However, it overlooks the relevance between the selected views and the input question, limiting its effectiveness in 3D-QA.
Table S2 presents the quantitative results on SQA (Ma et al., 2022), detailing performance across different question types: “What”, “Is”, “How”, “Can”, “Which”, and “Other”. Compared to state-of-the-art methods, LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ achieves the best performance on “What”, “How”, “Which”, and “Other” questions but shows a decline of 2.8% and 10.2% on “Is” and “Can” questions, respectively. This decline may be attributed to the zero-shot nature of LLAVA-OV (Li et al., 2024a), which maintains balanced performance across all question types. In contrast, other methods exhibit uneven performance, excelling in Is” and Can” questions due to dataset-specific adaptation while potentially underperforming in other categories. Furthermore, based on the same 2D LVLM, LLAVA-OV, cdViews consistently outperforms uniform sampling and image retrieval across all question types, demonstrating its effectiveness in selecting critical views for 3D-QA.
@@
GPT-3 (Brown et al., 2020) 3D 39.7 46.0 40.5 45.6 36.1 38.4 41.0
ScanQA (Azuma et al., 2022) 3D 28.6 65.0 47.3 66.3 43.9 42.9 45.3
SQA3D (Ma et al., 2022) 3D 33.5 66.1 42.4 69.5 43.0 46.4 47.2
3D-LLM (Hong et al., 2023) 3D 36.5 65.6 47.2 68.8 48.0 46.3 48.1
3D-VisTA (Zhu et al., 2023) 3D 34.8 63.3 45.4 69.8 47.2 48.1 48.5
SIG3D (Man et al., 2024a) 3D 35.6 67.2 48.5 71.4 49.1 45.8 52.6
PQ3D (Zhu et al., 2025) 3D+2D 37.1 61.3 44.5 60.9 47.0 45.1 47.1
BridgeQA (Mo & Liu, 2024) 3D+2D - - - - - - 52.9
@@
margin over the compared best $15.3\uparrow$ $-2.8\downarrow$ $5.5\uparrow$ $-10.2\downarrow$ $2.5\uparrow$ $6.3\uparrow$ $3.9\uparrow$
Table S2: Result comparisons with the state-of-the-art methods on the test set of the SQA (Ma et al., 2022). The best and second-best results are in bold and underlined. The decline in the “Is” and “Can” problems for LLAVA-OV with different view selections is attributed to the zero-shot nature of LLAVA-OV, which ensures balanced performance across all question types. In contrast, the compared methods exhibit uneven performance, excelling in “Is” and “Can” questions due to dataset-specific adaptation while potentially underperforming in other categories.
@@ Appendix C More Ablation Studies
This supplementary is for Section 5.2 of the main paper. The ablation studies are conducted on the validation set of ScanQA (Azuma et al., 2022), we evaluate the impact of different backbones, the effect of caption generation in viewAnnotator, and visualize the views within different distance thresholds, and visually compare the results of BridgeQA, LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$.
Ablation with Different Backbones. We evaluate the impact of different backbones by comparing LLAVA-NEXT (Li et al., 2024b) and LLAVA-OV (Li et al., 2024a), with results presented in Table S3. The results reveal two key insights: 1) View selection plays a crucial role in enhancing performance across models. Replacing uniform sampling with image retrieval improves performance by 1.1% on LLAVA-NEXT and 0.8% on LLAVA-OV, underscoring the importance of selecting informative views for 3D-QA. Our cdViews further amplifies these gains, achieving improvements of 3.6% and 1.8%, respectively, by effectively identifying more critical views. 2) cdViews demonstrates robustness and adaptability, consistently outperforming both baselines and delivering the highest performance gains across all evaluation metrics.
@@
Table S3: Ablation study results with different backbone models, LLAVA-Next (Li et al., 2024b) and LLAVA-OV (Li et al., 2024a). The best results are in bold. Subscripts indicate the relative improvement over the corresponding baseline, i.e., the 2D LVLM with uniform sampling.
@@
Effect of Finetuning LLaVA-OV in a Hybrid Method. To assess the feasibility and effectiveness of incorporating LLAVA-OV into a hybrid method, we implement a variant of BridgeQA—the strongest hybrid baseline in our main comparisons. Specifically, we retain the original BridgeQA architecture but replace its 2D vision-language module (BLIP (Li et al., 2022)) with LLAVA-OV. For a fair comparison, we also replace its top-1 image input with 9 views selected by our cdViews strategy, while keeping the full-scene point cloud features extracted by VoteNet (Qi et al., 2019).
During training, we adopt parameter-efficient tuning by updating only the last 2 of the 28 transformer layers in LLAVA-OV using LoRA (Hu et al., 2022). As shown in Table S5, the finetuned variant (BridgeQALLAVA-OV) achieves a +1.4% EM@1 improvement over the original BridgeQA baseline (28.4% vs. 27.0%), confirming the benefit of using a stronger 2D LVLM. Nonetheless, it still underperforms our cdViews, which achieves 30.1% EM@1. This experiment demonstrates that while hybrid pipelines can benefit from stronger LVLMs, they rely on complex architectures, 3D-specific modules, and computationally expensive fine-tuning. In contrast, our framework achieves superior performance by using a 2D-only LVLM in a zero-shot inference manner.
@@ Appendix D More Case Studies
Visualize Comparison of Different Methods and Their Visual Inputs. Figure S3 presents a visual comparison of the predicted answers and visual inputs of BridgeQA (Mo & Liu, 2024), LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$. BridgeQA relies on the top-1 image retrieval view combined with point clouds as input. However, relying on point clouds to provide the whole scene often results in answers that miss critical details. For instance, in the $4_{th}$ row, while the model correctly mentions the trash can on the floor, it overlooks surrounding objects like the toilet, which is crucial for providing a more informative answer. For LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, image retrieval-based view selection may miss the critical views required for accurate answers. As shown in the $3_{rd}$ rows, the retrieved views tend to be redundant or incomplete. In the $3_{rd}$ row, the selected views focus on the cabinet beneath the window but omit the view displaying books on top, which is essential for correctly answering the question. In contrast, LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ selects critical and diverse views, capturing essential context and delivering accurate, informative answers.
Figure S3: More Qualitative results for BridgeQA (Mo & Liu, 2024), LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and our final model LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$. Small marks Refer to caption, Refer to caption, and Refer to caption represents the selected views by each method. We can see that cdViews can capture critical and diverse views to answer the questions.

### v2

@@ 5.1 Comparisons with the State-of-the-Arts
Table 1 presents the quantitative results comparing 2D-only methods (uniform sampling, image retrieval, and cdViews) with other LLAVA-OV (li2024llava) with state-of-the-art 3D and hybrid methods. First, it is observed that 2D-only methods achieve superior performance, showing the advantage of applying 2D pre-trained models for 3D tasks. For example, compared to BridgeQA (mo2024bridging), our $\mathcal{F}_{\texttt{cdViews}}$ achieves significant improvements of 15.0% and 21.1% CIDEr on the two test sets of ScanQA. Second, among the 2D-only methods, cdViews outperforms the others. For example, $\mathcal{F}_{\texttt{cdViews}}$ outperforms $\mathcal{F}_{\text{retrieval}}$ by 4.0% and 2.7% CIDEr on both test sets of ScanQA. The reason is that the uniform sampling method ignores the question and the image retrieval method often fails to capture critical views or introduces redundancy views. In contrast, cdViews effectively identifies critical and diverse views for efficient 3D-QA. The qualitative comparison of selected views is shown in Figure 2 and Figure 6. More comparisons are provided in the Appendix Section B.
@@
+ $\mathcal{F}_{\text{uniform}}$ - - 28.3 17
+ $\mathcal{F}_{retrieval}$ - - 29.1 17
+ $\mathcal{F}_{retrieval}$ - 29.2 9
+ $\mathcal{F}_{\texttt{cdViews}}$ - 29.7 17
@@ 5.2 Ablation Study
In this section, we conduct an ablation study on the validation set of ScanQA (azuma2022scanqa), following (mo2024bridging). We study the impact of cdViews components and viewNMS thresholds. In addition, we particularly compare ours with the most related work: image-retrieval-based 3D-QA (mo2024bridging). More ablation studies are in Section C of the Appendix.
@@
Figure 7: The results of EM@1 using two configurations: optimal $k$ (blue) vs. fixed k=9k\!=\!9 (green). X-axis is the threshold $T$ of viewNMS. T=0T\!=\!0 means disabling viewNMS.
@@
Parameters 644644M 5.95.9M ($99.1\%$)
FLOPs 593.6593.6T 294.5294.5T ($50.4\%$)
Inference Time $2.8$s $1.2$s ($57.1\%$)
@@ Acknowledgments
This supplementary includes the details of view matching in viewAnnotation (Sec. A), more comparisons with the State-of-the-Arts (Section B), more ablation studies (Section C), including ablation of view selection methods with different 2D LVLM, effectiveness of caption generation in viewAnnotator, and more case studies (Section D).
@@ Appendix A More Details in View Matching
This supplementary is for Sec. 4.1 of the main paper. In our viewAnnotator, view matching classifies views as positive, negative, or uncertain. However, directly using a 2D LVLM with the prompt $Prompt_{M}$ as an instruction is unreliable, as the model lacks an explicit judgment criterion. To address this, we leverage its strong in-context learning ability (zhou2024visual) by providing a textual context example that guides the model through a structured reasoning process. Specifically, we incorporate a step-by-step system prompt in the View Matching process. As shown in Figure S1, the system prompt ensures that all key objects, attributes, and spatial relationships in the caption align with the image, reducing ambiguity and improving consistency. Uncertain views are explicitly excluded, enhancing the robustness of the annotation process. Additional examples of positive and negative views are shown in Figure S2.
@@
      ScanQA (azuma2022scanqa) 3D 20.3 29.5 32.4 61.7
      3D-LLM (hong20233d) 3D 20.5 39.3 35.7 69.4
      3D-VLP (Jin_2023_CVPR) 3D 21.7 30.5 34.5 67.0
      LL3DA (chen2024ll3da) 3D+2D - - 37.3 76.8
      BridgeQA (mo2024bridging) 3D+2D 27.0 - - -
      GPT-4O+CC (liu2024coarse) 2D - 35.4 42.6 87.0
@@
margin over the compared best $1\uparrow 3.3$$3\uparrow$ $4.2\uparrow$ $7.0\uparrow$
Table S1: Result comparisons with the state-of-the-art methods on the validation set of ScanQA (azuma2022scanqa). The best and second best results are in bold and underlined.
@@ Appendix B More Comparisons with the State-of-the-Art Methods
This supplementary is for Section 5.1 of the main paper. Table S1 presents the quantitative results comparing LLAVA-OV (li2024llava) with different view selection methods, including uniform sampling, image retrieval, and our cdViews, against state-of-the-art methods on the validation set of ScanQA. As shown, LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ outperforms these methods by clear margins. The last row of Table 1 highlights the performance gap between LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ and the best-performing baselines. Even compared to GPT-4O+CC (liu2024coarse), which leverages the powerful capabilities of GPT-4O (openai_gpt4o), LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ surpasses it by 7.0% CIDEr. GPT-4O+CC improves spatial understanding by adding object markers to track correspondences across uniformly sampled views. However, it overlooks the relevance between the selected views and the input question, limiting its effectiveness in 3D-QA.
Table S2 presents the quantitative results on SQA (ma2022sqa3d), detailing performance across different question types: “What”, “Is”, “How”, “Can”, “Which”, and “Other”. Compared to state-of-the-art methods, LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ achieves the best performance on “What”, “How”, “Which”, and “Other” questions but shows a decline of 2.8% and 10.2% on “Is” and “Can” questions, respectively. This decline may be attributed to the zero-shot nature of LLAVA-OV (li2024llava), which maintains balanced performance across all question types. In contrast, other methods exhibit uneven performance, excelling in Is” and Can” questions due to dataset-specific adaptation while potentially underperforming in other categories. Furthermore, based on the same 2D LVLM, LLAVA-OV, cdViews consistently outperforms uniform sampling and image retrieval across all question types, demonstrating its effectiveness in selecting critical views for 3D-QA.
@@
GPT-3 (brown2020language) 3D 39.7 46.0 40.5 45.6 36.1 38.4 41.0
ScanQA (azuma2022scanqa) 3D 28.6 65.0 47.3 66.3 43.9 42.9 45.3
SQA3D (ma2022sqa3d) 3D 33.5 66.1 42.4 69.5 43.0 46.4 47.2
3D-LLM (hong20233d) 3D 36.5 65.6 47.2 68.8 48.0 46.3 48.1
3D-VisTA (zhu20233dvista) 3D 34.8 63.3 45.4 69.8 47.2 48.1 48.5
SIG3D (man2024situational) 3D 35.6 67.2 48.5 71.4 49.1 45.8 52.6
PQ3D (zhu2025unifying) 3D+2D 37.1 61.3 44.5 60.9 47.0 45.1 47.1
BridgeQA (mo2024bridging) 3D+2D - - - - - - 52.9
@@
margin over the compared best $15.3\uparrow$ $2.8\downarrow$ $5.5\uparrow$ $10.2\downarrow$ $2.5\uparrow$ $6.3\uparrow$ $3.9\uparrow$
Table S2: Result comparisons with the state-of-the-art methods on the test set of the SQA (ma2022sqa3d). The best and second-best results are in bold and underlined. The decline in the “Is” and “Can” problems for LLAVA-OV with different view selections is attributed to the zero-shot nature of LLAVA-OV, which ensures balanced performance across all question types. In contrast, the compared methods exhibit uneven performance, excelling in “Is” and “Can” questions due to dataset-specific adaptation while potentially underperforming in other categories.
@@ Appendix C More Ablation Studies
This supplementary is for Section 5.2 of the main paper. The ablation studies are conducted on the validation set of ScanQA (azuma2022scanqa), we evaluate the impact of different backbones, the effect of caption generation in viewAnnotator, and visualize the views within different distance thresholds, and visually compare the results of BridgeQA, LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$.
Ablation with Different Backbones. We evaluate the impact of different backbones by comparing LLAVA-NEXT (li2024llavanext) and LLAVA-OV (li2024llava), with results presented in Table S3. The results reveal two key insights: 1) View selection plays a crucial role in enhancing performance across models. Replacing uniform sampling with image retrieval improves performance by 1.1% on LLAVA-NEXT and 0.8% on LLAVA-OV, underscoring the importance of selecting informative views for 3D-QA. Our cdViews further amplifies these gains, achieving improvements of 3.6% and 1.8%, respectively, by effectively identifying more critical views. 2) cdViews demonstrates robustness and adaptability, consistently outperforming both baselines and delivering the highest performance gains across all evaluation metrics.
@@
Table S3: Ablation study results with different backbone models, LLAVA-Next (li2024llavanext) and LLAVA-OV (li2024llava). The best results are in bold. Subscripts indicate the relative improvement over the corresponding baseline, i.e., the 2D LVLM with uniform sampling.
@@
Effect of Finetuning LLaVA-OV in a Hybrid Method. To assess the feasibility and effectiveness of incorporating LLAVA-OV into a hybrid method, we implement a variant of BridgeQA—the strongest hybrid baseline in our main comparisons. Specifically, we retain the original BridgeQA architecture but replace its 2D vision-language module (BLIP (li2022blip)) with LLAVA-OV. For a fair comparison, we also replace its top-1 image input with 9 views selected by our cdViews strategy, while keeping the full-scene point cloud features extracted by VoteNet (qi2019deep).
During training, we adopt parameter-efficient tuning by updating only the last 2 of the 28 transformer layers in LLAVA-OV using LoRA (hu2022lora). As shown in Table S5, the finetuned variant (BridgeQALLAVA-OV) achieves a +1.4% EM@1 improvement over the original BridgeQA baseline (28.4% vs. 27.0%), confirming the benefit of using a stronger 2D LVLM. Nonetheless, it still underperforms our cdViews, which achieves 30.1% EM@1. This experiment demonstrates that while hybrid pipelines can benefit from stronger LVLMs, they rely on complex architectures, 3D-specific modules, and computationally expensive fine-tuning. In contrast, our framework achieves superior performance by using a 2D-only LVLM in a zero-shot inference manner.
@@ Appendix D More Case Studies
Visualize Comparison of Different Methods and Their Visual Inputs. Figure S3 presents a visual comparison of the predicted answers and visual inputs of BridgeQA (mo2024bridging), LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$. BridgeQA relies on the top-1 image retrieval view combined with point clouds as input. However, relying on point clouds to provide the whole scene often results in answers that miss critical details. For instance, in the $4_{th}$ row, while the model correctly mentions the trash can on the floor, it overlooks surrounding objects like the toilet, which is crucial for providing a more informative answer. For LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, image retrieval-based view selection may miss the critical views required for accurate answers. As shown in the $3_{rd}$ rows, the retrieved views tend to be redundant or incomplete. In the $3_{rd}$ row, the selected views focus on the cabinet beneath the window but omit the view displaying books on top, which is essential for correctly answering the question. In contrast, LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$ selects critical and diverse views, capturing essential context and delivering accurate, informative answers.
Figure S3: More Qualitative results for BridgeQA (mo2024bridging), LLAVA-OV + $\mathcal{F}_{\text{retrieval}}$, and our final model LLAVA-OV + $\mathcal{F}_{\texttt{cdViews}}$. Small marks Refer to caption, Refer to caption, and Refer to caption represents the selected views by each method. We can see that cdViews can capture critical and diverse views to answer the questions.